# Bypassing the lattice BCS-BEC crossover in strongly correlated superconductors: resilient coherence from multiorbital physics

### v2

@@ Results
where $\hbar\underline{\bm{v}}=\nabla_{\bm{k}}\underline{h}(\bm{k})$ is the group velocity obtained from the one-electron Hamiltonian $\underline{h}(\bm{k})$ and the trace runs over the orbital indices of $\underline{\bm{v}}$ and $\underline{G}_{\bm{q}}$. Underlined quantities indicate matrices in orbital space, $N_{\bm{k}}$ is the number of momentum points and $e$ is the elementary charge. See Materials and Methods in the supplementary material for details on the DMFT-based implementation and sections S3, S4, and S5.B of the supplementary text for a derivation and discussion of Eq. (6).
The bottom panels of Fig. 2 show an example of our DMFT calculations which illustrates the $\bm{q}$-dependence of the OP amplitude and current density for different temperatures $T$. Throughout the paper, we choose $\bm{q}$ parallel to a reciprocal lattice vector $\bm{q}=q\bm{b}_{1}$. We find a monotonous suppression of the OP with increasing $q$. The supercurrent initially grows linearly with $q$, reaches its maximum $j_{\mathrm{dp}}$ and then collapses upon further increase of $q$. Thus, both $|\Psi_{q}|$ and $j_{q}$ behave qualitatively as expected from the GL description. For decreasing temperature, the point where the OP gets significantly suppressed moves towards larger momenta $q$ (smaller length scales $\xi$), while $j_{\mathrm{dp}}$ increases. We indicate the points where we extract $\xi$ and $j_{\mathrm{dp}}$ with gray circles connected by dashed lines.
@@
We apply FMP superconductivity to study a degenerate three-orbital model $H=H_{\mathrm{kin}}+H_{\mathrm{int}}$, where $H_{\mathrm{kin}}$ is the kinetic energy and the electron-electron interaction is described by a local Kanamori-Hubbard interaction (54)
@@
with total number $\hat{N}$, spin $\hat{\bm{S}}$, and angular momentum operator $\hat{\bm{L}}$. The independent interactions are the intraorbital Hubbard term $U$ and Hund exchange $J$, as we use the rotational SU(2)$\times$×SO(3) symmetric parametrization. This model is often discussed in the context of Hund’s metal physics relevant to, e.g., transition-metal oxides like ruthenates with partially filled $t_{2g}$ shells (54, 55). In the special case of negative exchange energy $J<0$, it has been introduced to explain superconductivity in A3C60 materials (56, 57, 58). In fullerides, exotic $s$-wave superconductivity exists in proximity to a Mott-insulating (MI) (28, 29, 59) and a Jahn-Teller metallic phase (29, 60, 61, 62). The influence of strong correlation effects and inverted Hund’s coupling were shown to be essential for the SC pairing (56, 57, 58, 63) utilizing orbital fluctuations (64) in a Suhl-Kondo mechanism (60).
The inversion of $J$ seems unusual from the standpoint of atomic physics, where it dictates the filling of atomic shells via Hund’s rules. In A3C60, a negative $J$ is induced by the electronic system coupling to intramolecular Jahn-Teller phonon modes (63, 65, 66). As a result, Hund’s rules are inverted such that states which minimize first spin $\bm{S}$ and then angular momentum $\bm{L}$ are energetically most favorable, see the second term of Eq. (7) for $J<0$.
We connect the model to A3C60 by using an ab initio derived model of the half-filled $t_{1u}$ bands for $H_{\mathrm{kin}}$ with the bandwidth $W$ as the unit of energy ($W\approx 0.3$ – $0.5$ eV for Cs to K based A3C60) (58, 67). For the interaction, we take the first principles’ estimates of fixed $J/W=-0.04$ and varying $U/W$ (56, 57, 58, 61) to emulate unit cell volumes as resulting from the size of the different alkali dopants. We show the temperature dependence of $|\Psi_{q=0}|$, $\xi$, and $\lambda_{\mathrm{L}}$ for different $U/W$ in Fig. 3.
Fig. 3: Order parameter, correlation length, and penetration depth in A3C60. The temperature dependence of (A) the zero-momentum order parameter $|\Psi_{0}|$, (B) the correlation length $\xi$, and (C) the London penetration depth $\lambda_{\mathrm{L}}$ are shown for different ratios $U/W$. The data were obtained for fixed $J/W=-0.04$ as estimated from ab initio data. Fits to extract critical temperatures $T_{\mathrm{c}}$ and zero-temperature values $\xi_{0}$ and $\lambda_{\mathrm{L},0}$ are shown by dashed lines. The region of uncertainty to fit $T_{\mathrm{c}}$ is indicated by shaded regions.
@@
Turning to the correlation length, we observe that, away from $T_{\mathrm{c}}$, $\xi$ is strongly reduced to only a few lattice constants ($a_{\mathrm{A}_{3}\mathrm{C}_{60}}\sim 14.2$ – 14.5 Å) by increasing $U$, i.e., pairing becomes very localized. At the same time, $\lambda_{\mathrm{L}}$ is enlarged. Hence, the condensate becomes much softer, i.e., there is a reduction of the SC stiffness $D_{\mathrm{s}}\propto\lambda_{\mathrm{L}}^{-2}$ upon increasing $U$. Fitting Eqs. (2) and (3) to our data, we obtain zero-temperature values of $\xi_{0}=3$ – 10 nm and $\lambda_{\mathrm{L},0}=80$ – 120 nm. Comparing our results with experimental values of $\xi_{0}\sim 2$ – 4.5 nm and $\lambda_{\mathrm{L}}\sim 200$ – 800 nm (6, 7, 30, 68, 69), we see an almost quantitative match for $\xi_{0}$ and a qualitative match for $\lambda_{\mathrm{L}}$. Both experiment and theory consistently classify A3C60 as type II superconductors ($\lambda_{\mathrm{L}}\gg\xi_{0}$(69, 30).
@@
We speculate that disorder and spontaneous orbital-symmetry breaking (62) in the vicinity of the Mott state could lead to a further reduction of $\xi_{0}$ as well as an increase of $\lambda_{\mathrm{L}}$ beyond what is found here for the pure system. This could bring our calculations with minimal $\xi_{0}=3$ nm closer to the experimental minimal coherence length of 2 nm revealed by measurements of large upper critical fields reaching up to a maximal $H_{\mathrm{c}2}=90$ T in Ref. 68 using $H_{\mathrm{c}2}=\Phi_{0}/(2\pi\xi_{0})$ with the flux quant $\Phi_{0}$.
@@
Fig. 4: Superconducting state of the A3C60 model in the ($U$, $J$)-interaction space. (A) Critical temperature $T_{\mathrm{c}}$, zero temperature penetration depth $\lambda_{\mathrm{L},0}$, inverse square root of quasiparticle weight $Z$, coherence length $\xi_{0}$, and the statistical weight $\rho_{\left|\upphi_{0}\right\rangle}^{N}$ of the local lowest energy states $\left|\upphi_{0}\right\rangle$ of the $N=2,3,4$ particle sectors obeying inverted Hund’s rules as a function of $U$ and $J$. Gray dots show original data points used for interpolation and the dashed line indicates a region where the proximity to the Mott state leads to a suppression of the superconducting state. There is no data point at the charge degeneracy line $U=2|J|$ in the lower left corner as marked by black ‘canceling’ lines. (B) Distribution of statistical weights $\rho_{\left|\upphi_{0}\right\rangle}^{N}$ at four different interaction values $U$ and $J$. (See section S7 in the supplementary text for a listing of the eigenstates $\left|\upphi_{0}\right\rangle$ and their respective local eigenenergies $E_{n}^{N}$). Red (blue) bars denote the density matrix weight of ground states in the $N=3$ ($N=2,4$) particle sector, the sum of which is plotted in the last panel of A. (C) Exemplary depiction of representative lowest inverted Hund’s rule eigenstates. A delocalized doublon (electron pair) fluctuates between different orbitals due to correlated pair hopping $J$.
@@
Circumvention of the lattice BCS-BEC crossover upon boosting inverted Hund’s coupling
@@
$\xi_{0}$, in contrast, depends strongly on $J$. By just slightly increasing $|J|$ above the ab initio estimate of $|J|$, the SC state becomes strongly localized with a short coherence length on the order of 2 – 3 $a_{\mathrm{A}_{3}\mathrm{C}_{60}}$. Remarkably, the small value of $\xi_{0}$ is independent of $U$ and thus the proximity to the MI phase. The localization of the Cooper pairs with $\xi_{0}$ on the order of the lattice constant is reminiscent of a BEC-type SC state. However, the dome-shaped behavior, characteristic of the lattice BCS-BEC crossover, with decreasing $T_{\mathrm{c}}$ in the strong coupling limit is notably absent here. Instead, $T_{\mathrm{c}}$ still grows inside the plateau of minimal $\xi_{0}$ when increasing the effective pairing strength proportional to $|J|$ for fixed $U/W$ (c.f. Fig. 5A). Only by diagonally traversing the ($U$, $J$) phase space, it is possible to suppress $T_{\mathrm{c}}$ inside the short $\xi_{0}$ plateau similar to Ref. 61.
@@
The reason for this circumvention of the lattice BCS-BEC phenomenology can be understood from an analysis of the local density matrix weights $\rho_{|\upphi_{n}\rangle}$, where $|\upphi_{n}\rangle$ refers to the eigenstates of the local Hamiltonian of our DMFT auxiliary impurity problem. We show $\rho_{|\upphi_{n}\rangle}$ of four different points in the interaction phase space in Fig. 4B. In the region of short $\xi_{0}$, the local density matrix is dominated by only eight states (red and blue bars) given by the “inverted Hund’s rule” ground states $|\upphi_{0}\rangle$ of the charge sectors with $N=2,3,4$ particles that are sketched in Fig. 4C. This can be seen in the last panel of Fig. 4A as the total weight of these eight states approaches one when entering the plateau of short $\xi_{0}$.
[…]
@@
Correspondingly, as the superconducting state at $-J/W>0.05$ relies on direct transitions between the local inverted Hund’s ground states from filling $N=3$ to $N=2$ and $N=4$, the local Hubbard repulsion $U$ has to fulfill two requirements for superconductivity with appreciable critical temperatures. First, significant occupation of the $N=2$ and 4 states at half-filling requires that $U$ is not too large. Otherwise, the system turns Mott insulating (around $U/W\gtrsim 1.6$ for enhanced $|J|$) and the SC phase stiffness is reduced upon increasing $U$ towards the Mott limit. At the same time, a significant amount of statistical weight of the $N=3$ states demands that $U$ must not be too small either. We find that $\Delta_{\mathrm{at}}>0$ and thus $U>2|J|$ is necessary for robust SC pairing. At $\Delta_{\mathrm{at}}<0$ there is a predominance of $N=2$ and $4$ states, which couple kinetically only in second-order processes and which are susceptible to charge disproportionation (70). Our analysis shows that a sweet spot for a robust SC state with high phase stiffness and large $T_{\mathrm{c}}$ exists for $|J|$ approaching $U/2$.
@@
Fig. 5: Crossover of superconducting properties between different interaction regimes in the A3C60 model. Critical temperature $T_{\mathrm{c}}$, coherence length $\xi_{0}$, and stiffness $D_{\mathrm{s}}\propto\lambda_{\mathrm{L},0}^{-2}$ when approaching the multiorbital strong coupling state as function of $|J|/W$ at fixed $U/W=0.8$ (A) and when approaching the Mott insulating state as function of $U/W$ at fixed $J/W=-0.04$ (B). To indicate a general trend, each quantity is normalized to its maximal value within the line cut. The phenomenology of the SC regimes resembles the BCS to lattice BEC crossover in the latter case (B), which is distinct from the multiorbital strong coupling case (A) (c.f. summary listed in Table 1). Note that the $T_{\mathrm{c}}$ at $U/W=2$ in (B) corresponds to the transition from a Mott insulating state instead of a metallic phase, as visually differentiated by a dashed connecting line.

### v4

@@ Results
where $\hbar\underline{\bm{v}}=\nabla_{\bm{k}}\underline{h}(\bm{k})$ is the group velocity obtained from the one-electron Hamiltonian $\underline{h}(\bm{k})$ and the trace runs over the orbital indices of $\underline{\bm{v}}$ and $\underline{G}_{\bm{q}}$. Underlined quantities indicate matrices in orbital space, $N_{\bm{k}}$ is the number of momentum points, and $e$ is the elementary charge. See Methods and Supplementary Note 4 for details on the DMFT-based implementation and Supplementary Notes 3 for a derivation and discussion of Eq. (6).
The bottom panels of Fig. 2 show an example of our DMFT calculations which illustrates the $\bm{q}$-dependence of the OP amplitude and current density for different temperatures $T$. Throughout the paper, we choose $\bm{q}$ parallel to a reciprocal lattice vector $\bm{q}=q\bm{b}_{1}$. We find a monotonous suppression of the OP with increasing $q$. The supercurrent initially grows linearly with $q$, reaches its maximum $j_{\mathrm{dp}}$ and then collapses upon further increase of $q$. Thus, both $|\Psi_{q}|$ and $j_{q}$ behave qualitatively as expected from the GL description. For decreasing temperature, the point where the OP gets significantly suppressed moves towards larger momenta $q$ (smaller length scales $\xi(T)$), while $j_{\mathrm{dp}}(T)$ increases. We indicate the points where we extract $\xi(T)$ and $j_{\mathrm{dp}}(T)$ with gray circles connected by dashed lines.
@@
We apply FMP superconductivity to study a degenerate three-orbital model $H=H_{\mathrm{kin}}+H_{\mathrm{int}}$, where $H_{\mathrm{kin}}$ is the kinetic energy and the electron-electron interaction is described by a local Kanamori–Hubbard interaction 69
@@
with total number $\hat{N}$, spin $\hat{\bm{S}}$, and angular momentum operator $\hat{\bm{L}}$. The independent interactions are the intraorbital Hubbard term $U$ and Hund exchange $J$, as we use the SU(2)$\times$×SO(3) symmetric parametrization. This model is often discussed in the context of Hund’s metal physics relevant to, e.g., transition-metal oxides like ruthenates with partially filled $t_{2g}$ shells 69, 70. In the special case of negative exchange energy $J<0$, it has been introduced to explain superconductivity in A3C60 materials 65, 66, 67, 68. In fullerides, exotic $s$-wave superconductivity exists in proximity to a Mott-insulating (MI) 35, 36, 71 and a Jahn–Teller metallic phase 36, 72, 73, 74. The influence of strong correlation effects and inverted Hund’s coupling were shown to be essential for the SC pairing 65, 66, 67, 68, 75 utilizing orbital fluctuations 76 in a Suhl–Kondo mechanism 72. In the weakly correlated regime of $U\to 0$ and small $J$, the system follows BCS phenomenology 66.
The inversion of $J$ seems unusual from the standpoint of atomic physics, where it dictates the filling of atomic shells via Hund’s rules. In A3C60, a negative $J$ is induced by the electronic system coupling to intramolecular Jahn–Teller phonon modes 75, 77, 78, 68. As a result, Hund’s rules are inverted such that states which minimize first spin $\bm{S}$ and then angular momentum $\bm{L}$ are energetically most favorable, see the second term of Eq. (7) for $J<0$.
We connect the model to A3C60 by using an ab initio derived model for the kinetic energy 79
$\begin{split}H_{\mathrm{kin}}&=\sum_{ij}\sum_{\alpha\gamma\sigma}t_{\alpha\gamma}(\bm{R}_{ij})c^{\dagger}_{i\alpha\sigma}c_{j\gamma\sigma}\end{split}$ (8)
where $t_{\alpha\gamma}(\bm{R}_{ij})$ is the hopping amplitude between half-filled $t_{1u}$ orbitals ($\alpha,\gamma=1,2,3$) of C60 molecules on sites connected by lattice vector $\bm{R}_{ij}$. We take the bandwidth $W$ as the unit of energy ($W\approx 0.3$ – $0.5$ eV for Cs to K based A3C6068, 79, see Supplemental Note 4-D for further details. For the interaction, we take the first principles’ estimates of fixed $J/W=-0.04$ and varying $U/W$ 65, 66, 67, 68, 73 to emulate unit cell volumes as resulting from the size of different alkali dopants. We solve this Hamiltonian using DMFT, which explicitly takes into account superconducting order. Through the momentum dependence of $|\Psi_{\bm{q}}(T)|$ and ${\bm{j}}_{\bm{q}}(T)$, $\xi(T)$ and $\lambda_{\mathrm{L}}(T)$ can be extracted as discussed in the previous subsection. We show the derived temperature dependence of $|\Psi_{q=0}(T)|$, $\xi(T)$, and $\lambda_{\mathrm{L}}(T)$ for different $U/W$ in Fig. 3.
Fig. 3: Order parameter, correlation length, and penetration depth in A3C60. The temperature dependence of (a) the zero-momentum order parameter $|\Psi_{0}|$, (b) the correlation length $\xi$, and (c) the London penetration depth $\lambda_{\mathrm{L}}$ are shown for different ratios $U/W$. The data were obtained for fixed $J/W=-0.04$ as estimated from ab initio data. Length scales are given in units of the lattice constant $a$. Fits to extract critical temperatures $T_{\mathrm{c}}$ and zero-temperature values $\xi_{0}$ and $\lambda_{\mathrm{L},0}$ are shown by dashed lines. The region of uncertainty to fit $T_{\mathrm{c}}$ is indicated by shaded regions.
@@
Turning to the correlation length, we observe that, away from $T_{\mathrm{c}}$, $\xi(T)$ is strongly reduced to only a few lattice constants ($a\sim 14.2$ – 14.5 Å) by increasing $U$, i.e., pairing becomes very localized. At the same time, $\lambda_{\mathrm{L}}(T)$ is enlarged. Hence, the condensate becomes much softer as there is a reduction of the SC stiffness $D_{\mathrm{s}}\propto\lambda_{\mathrm{L}}^{-2}$ upon increasing $U$. Fitting Eqs. (2) and (3) to our data, we obtain zero-temperature values of $\xi_{0}=3$ – 10 nm and $\lambda_{\mathrm{L},0}=80$ – 120 nm. Comparing our results with experimental values of $\xi_{0}\sim 2$ – 4.5 nm and $\lambda_{\mathrm{L}}\sim 200$ – 800 nm 6, 7, 37, 80, 81, we see an almost quantitative match for $\xi_{0}$ and a qualitative match for $\lambda_{\mathrm{L}}$. Both experiment and theory consistently classify A3C60 as type II superconductors ($\lambda_{\mathrm{L}}\gg\xi_{0}$81, 37.
@@
We speculate that disorder and spontaneous orbital-symmetry breaking 74 in the vicinity of the Mott state could lead to a further reduction of $\xi_{0}$ as well as an increase of $\lambda_{\mathrm{L}}$ beyond what is found here for the pure system. This could bring our calculations with minimal $\xi_{0}=3$ nm closer to the experimental minimal coherence length of 2 nm revealed by measurements of large upper critical fields reaching up to a maximal $H_{\mathrm{c}2}=90$ T in Ref. 80 using $H_{\mathrm{c}2}=\Phi_{0}/(2\pi\xi_{0})$ with the flux quant $\Phi_{0}$.
@@
Fig. 4: Superconducting state of the A3C60 model in the ($U$, $J$)-interaction space. (a) Critical temperature $T_{\mathrm{c}}$, zero temperature penetration depth $\lambda_{\mathrm{L},0}$, inverse square root of quasiparticle weight $Z$, coherence length $\xi_{0}$, and the statistical weight $\rho_{\left|\upphi_{0}\right\rangle}^{N}$ of the local lowest energy states $\left|\upphi_{0}\right\rangle$ of the $N=2,3,4$ particle sectors obeying inverted Hund’s rules as a function of $U$ and $J$. Gray dots show original data points used for interpolation and the dashed line indicates a region where the proximity to the Mott state leads to a suppression of the superconducting state. There is no data point at the charge degeneracy line $U=2|J|$ in the lower left corner as marked by black ‘canceling’ lines. (b) Distribution of statistical weights $\rho_{\left|\upphi_{n}\right\rangle}^{N}$ at four different interaction values $U$ and $J$. (See Supplementary Note 7 for a listing of the eigenstates $\left|\upphi_{n}\right\rangle$ and their respective local eigenenergies $E_{n}^{N}$). Red (blue) bars denote the density matrix weight of ground states in the $N=3$ ($N=2,4$) particle sector, the sum of which is plotted in the last panel of a. (c) Exemplary depiction of representative lowest inverted Hund’s rule eigenstates. A delocalized doublon (electron pair) fluctuates between different orbitals due to correlated pair hopping $J$.
@@
Circumvention of the lattice BCS–BEC crossover upon boosting inverted Hund’s coupling
@@
$\xi_{0}$, in contrast, depends strongly on $J$. By just slightly increasing $|J|$ above the ab initio estimate of $|J|$, the SC state becomes strongly localized with a short coherence length on the order of 2 – 3 $a$. Remarkably, the small value of $\xi_{0}$ is independent of $U$ and thus the proximity to the MI phase. The localization of the condensate with $\xi_{0}$ on the order of the lattice constant is reminiscent of a crossover to the BEC-type SC state. However, the dome-shaped behavior, characteristic of the lattice BCS–BEC crossover 27, 29, with decreasing $T_{\mathrm{c}}$ in the strong coupling limit is notably absent here. Instead, $T_{\mathrm{c}}$ still grows inside the plateau of minimal $\xi_{0}$ when increasing the effective pairing strength proportional to $|J|$ for fixed $U/W$ (c.f. Fig. 5a). Only by diagonally traversing the ($U$, $J$) phase space, it is possible to suppress $T_{\mathrm{c}}$ inside the short $\xi_{0}$ plateau with a dome structure as shown, e.g., in Ref. 73.
@@
The reason for this circumvention of the lattice BCS–BEC phenomenology can be understood from an analysis of the local density matrix weights $\rho_{|\upphi_{n}\rangle}$, where $|\upphi_{n}\rangle$ refers to the eigenstates of the local Hamiltonian of our DMFT auxiliary impurity problem. We show $\rho_{|\upphi_{n}\rangle}$ of four different points in the interaction phase space in Fig. 4b. In the region of short $\xi_{0}$, the local density matrix is dominated by only eight states (red and blue bars) given by the “inverted Hund’s rule” ground states $|\upphi_{0}\rangle$ of the charge sectors with $N=2,3,4$ particles that are sketched in Fig. 4c. This can be seen in the last panel of Fig. 4a as the total weight of these eight states approaches one when entering the plateau of short $\xi_{0}$.
By increasing $|J|$, the system is driven into a strong coupling phase where local singlets are formed as Cooper pair precursors 77 while electronic hopping is not inhibited. On the contrary, hopping between the $N=2,3,4$ ground states is even facilitated via large negative $J$. It does so by affecting two different energy scales (c.f. Supplementary Note 7): Enhancing $J<0$ reduces the atomic gap $\Delta_{\mathrm{at}}=E_{0}^{N=4}+E_{0}^{N=2}-2E_{0}^{N=3}=U-2|J|$ 69 relevant to charge excitations and thereby supports hopping. A higher negative Hund’s exchange simultaneously increases the energy $\Delta E=2|J|$ necessary to break up the orbital singlets within a fixed charge sector. As a result, unpaired electrons in the $N=3$ state become more itinerant while the local Cooper pair binding strength increases. Since the hopping is not reduced, the SC stiffness is not compromised by larger $|J|$. This two-faced role or Janus effect of negative Hund’s exchange, that localizes Cooper pairs but delocalizes electrons, can be understood as a competition of a Mott and a charge disproportionated insulator giving way for a mixed-valence metallic state in between 82.
@@
Correspondingly, as the superconducting state at $-J/W>0.05$ relies on direct transitions between the local inverted Hund’s ground states from filling $N=3$ to $N=2$ and $N=4$, the local Hubbard repulsion $U$ has to fulfill two requirements for superconductivity with appreciable critical temperatures. First, significant occupation of the $N=2$ and 4 states at half-filling requires that $U$ is not too large. Otherwise, the system turns Mott insulating (around $U/W\gtrsim 1.6$ for enhanced $|J|$) and the SC phase stiffness is reduced upon increasing $U$ towards the Mott limit. At the same time, a significant amount of statistical weight of the $N=3$ states demands that $U$ must not be too small either. We find that $\Delta_{\mathrm{at}}>0$ and thus $U>2|J|$ is necessary for robust SC pairing. At $\Delta_{\mathrm{at}}<0$ there is a predominance of $N=2$ and $4$ states, which couple kinetically only in second-order processes and which are susceptible to charge disproportionation 82. Our analysis shows that a sweet spot for a robust SC state with high phase stiffness and large $T_{\mathrm{c}}$ exists for $|J|$ approaching $U/2$.
@@
Fig. 5: Crossover of superconducting properties between different interaction regimes in the A3C60 model. Critical temperature $T_{\mathrm{c}}$, coherence length $\xi_{0}$, and stiffness $D_{\mathrm{s}}\propto\lambda_{\mathrm{L},0}^{-2}$ when approaching the multiorbital strong coupling state as function of $|J|/W$ at fixed $U/W=0.8$ (a) and when approaching the Mott insulating state as function of $U/W$ at fixed $J/W=-0.04$ (b). To indicate a general trend, each quantity is normalized to its maximal value within the line cut. The phenomenology of the SC regimes resembles the BCS to lattice BEC crossover when approaching the Mott insulator (b), which is distinct from the multiorbital strong coupling case (a) (c.f. summary listed in Table 1). Note that the $T_{\mathrm{c}}$ at $U/W=2$ in (b) corresponds to the transition from a Mott insulating state instead of a metallic phase, as visually differentiated by a dashed connecting line.